# Reward-Guided Speculative Decoding for Efficient LLM Reasoning

### v2

@@ 1 Introduction
These advancements result in high computational and economic costs. While training is resource-intensive, inference at scale is even costlier, requiring vast computational infrastructure and energy to serve billions of queries (Patterson et al., 2021). The exponential growth in inference costs makes it a key challenge for large model deployment, highlighting the need for efficient techniques (Frantar et al., 2022; Lin et al., 2024; Xu et al., 2023; Sun et al., 2023; Zhang et al., 2023b; Li et al., 2024a; Xu et al., 2024; Liao & Monz, 2024; Liao et al., 2024a) to reduce energy use and ensure scalability.
Specifically, Sequential token generation in large LLMs incurs significantly higher computational costs compared to smaller models. This increased latency can hinder their deployment in real-time or high-throughput applications. To address this issue, parallel decoding techniques, such as speculative decoding, have emerged as effective solutions (Leviathan et al., 2023). Speculative decoding operates by leveraging a smaller, lightweight model to generate candidate outputs, which are then validated and refined by the larger model. This approach significantly reduces the number of decoding tokens required by the large model, thereby accelerating the overall process. The smaller model serves as a guide, proposing sequences that the larger model can confirm or adjust, leading to faster inference times without compromising quality. Furthermore, speculative decoding ensures efficiency by maintaining high-quality outputs through careful calibration of the smaller model. By aligning the smaller model’s predictions with the larger model’s capabilities, this method minimizes discrepancies and enhances reliability during inference.
@@ 2.2 Acceptance Criterion $\mathcal{A}_{\omega}$ and Weighting Function
Weighting Function Definition
$\omega(r)=p$ Constant $p\in(0,1)$
$\omega(r)=\mathbf{1}(r\geq\delta)$ 1 if $r\geq\delta$, else 0
$\omega(r)=\min(1,\max(0,r))$ Clipping $r$ within [0,1]
$\omega(r)=\max(0,\frac{r}{1+r})$ Sigmoidal transformation
$\omega(r)=\frac{1}{1+e^{-\alpha(r-\delta)}}$ Logistic function
@@ 3 Empirical Results
Datasets. We evaluate our method on a diverse set of reasoning tasks, including GSM8K (Cobbe et al., 2021b), MATH500 (Hendrycks et al., 2021), MMLU STEM (Hendrycks et al., 2020), OlympiadBench (He et al., 2024), GaoKao-2023-En (Liao et al., 2024b), GPQA (Rein et al., 2023), and Minerva Math (Lewkowycz et al., 2022).
@@ 3.1 Reasoning Benchmarks
Additionally, a larger PRM (7B) slightly enhances performance compared to a smaller PRM (1.5B), especially on complex datasets like GPQA and MATH500, where the increased reasoning capacity of a larger PRM proves beneficial. Results with general models, such as Qwen2.5-Instruct and Llama-3.1-Instruct, align with those from math-specific models, validating RSD’s robustness and generalizability.
@@
Beam Search bs=4 78.2 88.9 33.5
Beam Search bs=8 78.2 88.4 32.4
@@ 4 Discussion
Specialized PRM. In our experiments, we rely on an open-source general PRM. However, training or fine-tuning a specialized PRM that is closely aligned with the draft model could further enhance performance. Such a PRM would better recognize high-quality reasoning steps generated by the draft model, thereby increasing the acceptance rate. Future work could explore specialized PRM training or fine-tuning, leveraging targeted datasets and chain-of-thought annotations. This more tightly coupled feedback loop is especially promising in applications like mathematics, coding, or scientific discovery, where domain knowledge and consistent reasoning styles can amplify the benefits of reward-guided speculative decoding.
@@ 5 Related Work
Reward Models on Reasoning. Reward models play a crucial role in selecting correct reasoning trajectories during both training (Chen et al., 2024a; Wang et al., 2024) and inference (Brown et al., 2024). Outcome Reward Models (ORMs) (Yu et al., 2023; Dong et al., 2024) are trained exclusively on the model’s final output, whereas Process Reward Models (PRMs) (Lightman et al., 2023) rely on step-level annotations, providing dense and granular reward signals at each reasoning step. Scaling test-time compute (Snell et al., 2024; OpenAI, 2024) has gained significant traction with the advancement of reward models. Techniques like Best-of-$N$ (Brown et al., 2024; Cobbe et al., 2021a; Dong et al., 2023) leverage ORMs to select the sample with the highest reward from N candidates. Building on this, tree search methods have been introduced, enabling per-step predictions rather than relying solely on the final answer (Chen et al., 2024a; Qi et al., 2024; Yao et al., 2024). These methods are enhanced by process feedback, such as Process Reward Models (PRMs). We propose a novel application of PRMs to accelerate reasoning during inference.
6 Conclusion

### v3

@@ 1 Introduction
These advancements result in high computational and economic costs. While training is resource-intensive, inference at scale is even costlier, requiring vast computational infrastructure and energy to serve billions of queries (Patterson et al., 2021). The exponential growth in inference costs makes it a key challenge for large model deployment, highlighting the need for efficient techniques (Frantar et al., 2022; Lin et al., 2024; Xu et al., 2023; Sun et al., 2023; Zhang et al., 2023b; Li et al., 2024a; Xu et al., 2024; Liao & Monz, 2024) to reduce energy use and ensure scalability.
Specifically, sequential token generation in large LLMs incurs significantly higher computational costs compared to smaller models. This increased latency can hinder their deployment in real-time or high-throughput applications. To address this issue, parallel decoding techniques, such as speculative decoding, have emerged as effective solutions (Leviathan et al., 2023). Speculative decoding operates by leveraging a smaller, lightweight model to generate candidate outputs, which are then validated and refined by the larger model. This approach significantly reduces the number of decoding tokens required by the larger model, thereby accelerating the overall process. The smaller model serves as a guide, proposing sequences that the larger model can confirm or adjust, leading to faster inference without compromising quality. Furthermore, speculative decoding ensures efficiency by maintaining high-quality outputs through careful calibration of the smaller model. By aligning the smaller model’s predictions with the larger model’s capabilities, this method minimizes discrepancies and enhances reliability during inference.
@@ 2.2 Acceptance Criterion $\mathcal{A}_{\omega}$ and Weighting Function
Weighting Function Definition
$\omega(r)=p$ Constant $p\in(0,1)$
$\omega(r)=\mathbf{1}(r\geq\delta)$ 1 if $r\geq\delta$, else 0
$\omega(r)=\min(1,\max(0,r))$ Clipping $r$ within [0,1]
$\omega(r)=\max(0,\frac{r}{1+r})$ Sigmoidal transformation
$\omega(r)=\frac{1}{1+e^{-\alpha(r-\delta)}}$ Logistic function
@@ 3 Empirical Results
Datasets. We evaluate our method on a diverse set of reasoning tasks, including GSM8K (Cobbe et al., 2021b), MATH500 (Hendrycks et al., 2021), MMLU STEM (Hendrycks et al., 2020), OlympiadBench (He et al., 2024), GaoKao-2023-En (Liao et al., 2024), GPQA (Rein et al., 2023), and Minerva Math (Lewkowycz et al., 2022).
@@ 3.1 Reasoning Benchmarks
Additionally, a larger PRM (7B) slightly enhances performance compared to a smaller PRM (1.5B), especially on complex datasets like GPQA and MATH500, where the increased reasoning capacity of a larger PRM proves beneficial. Results with general models, such as Qwen2.5-Instruct and Llama-3.1-Instruct, are even improved more, validating RSD’s robustness and generalizability.
@@
Beam Search beam size $=4$ 78.2 88.9 33.5
Beam Search beam size $=8$ 78.2 88.4 32.4
@@ 4 Discussion
Table 5: Accuracy on reasoning benchmarks with different PRMs. RSD is robust to the choice of PRM. $\delta=0.7$ works well universally.
Method PRM Setting MATH500 GSM8K GaoKao Olympiad GPQA MMLU Avg.
2023 En Bench Diamond STEM
Draft: Qwen2.5-Math-Instruct-1.5B, Target: Qwen2.5-Math-Instruct-7B
Single Target Model - - 83.2 95.7 66.8 41.2 32.8 71.8 65.3
SD - - 83.4 95.6 67.3 40.6 28.8 72.0 64.6 (-0.7)
RSD Skywork-o1-Open-PRM-1.5B $\delta=0.7$ 82.6 94.5 68.8 39.6 38.4 71.4 65.9 (+0.6)
RSD Skywork-o1-Open-PRM-7B $\delta=0.7$ 84.6 95.5 68.3 42.1 33.8 72.3 66.1 (+0.8)
RSD Qwen2.5-Math-PRM-7B $\delta=0.7$ 83.2 95.5 67.3 41.3 35.3 72.5 65.9 (+0.6)
RSD Qwen2.5-Math-PRM-72B $\delta=0.7$ 84.0 95.7 68.1 42.0 33.3 75.6 66.5 (+1.2)
Robustness to PRMs. To investigate RSD’s robustness to the choice of PRM, we include two more strong PRMs, i.e. Qwen2.5-Math-PRM-7B and 72B (Zhang et al., 2025) in Table 5. RSD equipped with different PRMs consistently outperforms SD, with more gains from a larger PRM.
Table 6: Results on a general-domain benchmark, AlpacaEval. The PRM used here is an outcome reward model, Skywork-Reward-Llama-3.1-8B-v0.2.
Method Win Rate (%)
Draft Model only (Llama-3.2-1B-Instruct) 7.09
Target Model only (Llama-3.1-8B-Instruct) 24.47
RSD 18.85
General-Domain Task. An important component of RSD is PRM. To the best of our knowledge, there is not yet a PRM for general-domain generation. However, there are many outcome reward models (ORMs) for open-ended generation. Could we use an ORM instead of PRM in RSD?
[…]
As shown in the In Table 6, even with an ORM instead of a PRM, RSD achieves a significantly better win rate than the draft model, showing RSD’s robustness across different tasks. Among all generated tokens, 65% tokens are generated by the draft model only without any intervention of the target model. We believe that a general-domain PRM and dedicated tuning of $\delta$ could further boost the performance.
@@
Specialized PRM. In our experiments, we rely on an open-source general PRM. However, training or fine-tuning a specialized PRM that is closely aligned with the draft model could further enhance performance. Such a PRM would better recognize high-quality reasoning steps generated by the draft model, thereby increasing the acceptance rate. Future work could explore specialized PRM training or fine-tuning.
@@ 5 Related Work
Reward Models on Reasoning. Reward models play a crucial role in selecting correct reasoning trajectories during training (Chen et al., 2024a; Wang et al., 2024; Zhou et al., 2025) and inference (Brown et al., 2024). Outcome Reward Models (ORMs) (Yu et al., 2023; Dong et al., 2024) are trained exclusively on the model’s final output, whereas process reward models (PRMs) (Lightman et al., 2023) rely on step-level annotations, providing dense and granular reward signals at each reasoning step. Scaling test-time compute (Snell et al., 2024; OpenAI, 2024) has gained significant traction with the advancement of reward models. Techniques like Best-of-$N$ (Brown et al., 2024; Cobbe et al., 2021a; Dong et al., 2023) leverage ORMs to select the sample with the highest reward from N candidates. Building on this, tree search methods have been introduced, enabling per-step predictions rather than relying solely on the final answer (Chen et al., 2024a; Qi et al., 2024; Yao et al., 2024). These methods are enhanced by process feedback, such as process reward models (PRMs). We propose a novel application of PRMs to accelerate reasoning during inference.
6 Conclusion.
@@
Acknowledgements
This research was partly supported by the Netherlands Organization for Scientific Research (NWO) under project number VI.C.192.080.